Required: 1. What is last years margin? Margin

2. What is last years turnover? (Round your answer to 1 decimal place.) urnover

3. What is last years return on investment (ROI)? ROI

4. What is the margin related to this years investment opportunity? Margin

5. What is the turnover related to this years investment opportunity? (Round your answer to 1 decimal place.) urnover

6. What is the ROl related to this years investment opportunity? ROI

11. What is last years residual income? esidual income

12. What is the residual income of this years investment opportunity? Residual income

Required information The following information applies to the questions displayed below. Westerville Company reported the following results from last year's operations: Sales $ 2,000,000 Variable expenses 640,000 Contribution margin Fixed expenses 1,360,000 860,000 $ 500,000 Net operating income Average operating assets 1,250,000 At the beginning of this year, the company has a $250,000 investment opportunity with the following cost and revenue characteristics Sales $ 400,000 70 of sales Contribution margin ratio Fixed expenses $220,000 The company's minimum required rate of return is 10%.
Required: 1. What is last year's margin? Margin
2. What is last year's turnover? (Round your answer to 1 decimal place.) urnover
3. What is last year's return on investment (ROI)? ROI
4. What is the margin related to this year's investment opportunity? Margin
5. What is the turnover related to this year's investment opportunity? (Round your answer to 1 decimal place.) urnover
6. What is the ROl related to this year's investment opportunity? ROI
7. If the company pursues the investment opportunity and otherwise performs the same as last year, what margin will it earn this year? Round your percentage answer to 1 decimal place (i.e .1234 should be entered as 12.3)) argin
8. If the company pursues the investment opportunity and otherwise performs the same as last year, what turnover will it earn this year? (Round your answer to 2 decimal places.) urnover
9. If the company pursues the investment opportunity and otherwise performs the same as last year, what ROl will it earn this year? (Round your percentage answer to 1 decimal place (i.e., 0.1234 should be considered as 12.3%.))
11. What is last year's residual income? esidual income
12. What is the residual income of this year's investment opportunity? Residual income
13. If the company pursues the investment opportunity and otherwise performs the same as last year, what residual income will it earn this year? Residual income
Answer #1

1) Last year Margin = 500,000/2,000,000

=25%

2)Last Years Turnover = 2,000,000/1,250,000

=1.6

3)ROI = 25% x 1.6 = 40%

4) Margin = 60,000/400,000 =15%

5)Turnover = 400,000/250,000 = 1.6

6)15% x 1.6 = 24%

7)

Sales $ 2,000,000.00 $ 400,000.00 $ 2,400,000.00
Variable Expense $      640,000.00 $ 120,000.00 $      760,000.00
Contribution Margin $ 1,360,000.00 $ 280,000.00 $ 1,640,000.00
Fixed Expense $      860,000.00 $ 220,000.00 $ 1,080,000.00
Net Operating income $      500,000.00 $    60,000.00 $      560,000.00

Margin = 560,000/2,400,000

=23.33%

8)Turnover = 2,400,000 /1,500,000

=1.6

9)23.33% x 1.6 = 37.33%

11)500,000 - (1,250,000 x 10%)

=375,000

12)60,000 - (250,000 x 10%)

=35,000

13)560,000 - (1,500,000 x 10%)

=410,000